# On the backward stability of s-step GMRES

### v2

@@ Abstract
Communication, i.e., data movement, is a critical bottleneck for the performance of classical Krylov subspace method solvers on modern computer architectures. Variants of these methods which avoid communication have been introduced, which, while equivalent in exact arithmetic, can be unstable in finite precision. In this work, we address the backward stability of $s$-step GMRES, also known as communication-avoiding GMRES. Compared to the “modular framework” proposed in [A. Buttari, N. J. Higham, T. Mary, & B. Vieublé. Preprint in 2024.], we present an improved framework for simplifying the analysis of $s$-step GMRES, which includes standard GMRES ($s=1$) as a special case, by isolating the effects of rounding errors in the QR factorization and the solution of the least squares problem. The key advantage of this new framework is that it is evident how the orthogonalization method affects the backward error, and it is not necessary to re-evaluate anything other than the orthogonalization itself when modifying the orthogonalization used in GMRES. Using this framework, we analyze $s$-step GMRES with popular block orthogonalization methods: block modified Gram–Schmidt and reorthogonalized block classical Gram–Schmidt algorithms.
@@ 1 Introduction
The standard GMRES algorithm creates and orthogonalizes the Krylov basis one vector at a time, predominantly making use of BLAS 2 operations. On modern computer architectures, the performance of classical iterative solvers is heavily dominated by communication, i.e., data movement and synchronization. This motivated the introduction of $s$-step (also called communication-avoiding) variants of the GMRES algorithm, which can reduce asymptotic latency costs by a factor of $s$ and can take advantage of BLAS 3 operations. In [14, Section 3.6], it is demonstrated that CA-GMRES provides a speedup ranging from 1.31×1.31$\times$1.31 × to 4.3×4.3$\times$4.3 × compared to the standard GMRES on an Intel platform with 8 cores. The authors in [22] show that CA-GMRES can achieve speedups of up to 2×2$\times$2 × over the standard GMRES on one or multiple GPUs. Particularly, the orthogonalization phase in CA-GMRES can reach up to 4.16×4.16$\times$4.16 × speedup compared to the same phase in the standard GMRES when $s=10$ is used.
@@ 3 An abstract framework for backward stability of $s$-step GMRES
In GMRES, the orthogonalization method employed in Step 2 can be viewed as being applied to the matrix $\begin{bmatrix}\phi\widehat{r}&\widehat{W}_{1:(i+1)s}\end{bmatrix}$ for any positive scalar $\phi$. We assume the backward stability of the block orthogonalization method used in Step 2; according to [10], most commonly-used block orthogonalization schemes satisfy this property, i.e., for any $\phi>0$,
@@
From standard rounding error analysis and [13, Equation (3.12)], we have $\varepsilon_{r}$, $\varepsilon_{AZ}$, $\varepsilon_{mv}=O(\bm{u})$. In [5, Equation (A.7)], the authors have given $\varepsilon_{ls}=O(\bm{u})$, which comes from applying Givens QR factorization to the upper Hessenberg matrix $\hat{R}_{1:p+1,2:p+1}$ and solving the triangular system as depicted in Step 3.
@@
According to [19, Theorem 2.4], the residual of the least squares problem in Step 3, i.e., $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$, can be bounded by $\sigma_{\min}\bigg{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\bigg{)}$ and $\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$. Therefore, establishing a bound for the residual $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$ requires determining the relationship between the result and the assumptions (13) and (14), which amounts to bounding
@@
Lemma 1 gives bounds on the smallest and largest singular values of the matrices $\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}$ and $\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}$, which will be required for the proof of Lemma 3. We are now prepared to bound the residual of the least squares problem in Step 3, i.e., $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$, using [19, Theorem 2.4].
@@
We follow the approach from [5, 19] and employ [19, Theorem 2.4] to estimate the residual of the least squares problem. Let $\tilde{R}_{1:p+1}=\widehat{R}_{1:p+1}+\Delta R_{1:p+1}$. Then using [19, Theorem 2.4], the residual of the least squares problem can be estimated as, for any $\phi>0$,
@@ 4.1 Backward error of $s$-step GMRES with block Householder QR and block modified Gram–Schmidt (BMGS)
From [13, Theorem 19.4] and [23], it is easy to see that
@@
Then we consider the modified Gram–Schmidt (MGS) algorithm. [13, Theorem 19.13] and [4] showed (55) can be satisfied by establishing the equivalence between the MGS algorithm and the Householder QR algorithm. Similarly, it has been proved in [2] that applying the BMGS algorithm with Householder QR as the intra-block orthogonalization routine to $X$ is equivalent to applying the block Householder QR algorithm to $\begin{bmatrix}0_{n}\\ X\end{bmatrix}$ both mathematically and numerically. Then we obtain that (55) holds also for the BMGS algorithm.
@@
holds, as established in [16, Theorem 4.1]. Moreover, the contrapositive indicates that (13) is satisfied with $\omega_{\kappa(r,W)}=O(\bm{u})$ when there exists $p\leq n$ such that $\bigl{\|}\widehat{V}_{1:p+1}^{\top}\widehat{V}_{1:p+1}-I\bigr{\rVert}_{\mathsf{F}}>1/2$ but $\bigl{\|}\widehat{V}_{1:p}^{\top}\widehat{V}_{1:p}-I\bigr{\rVert}_{\mathsf{F}}\leq 1/2$.
@@ 4.2 Backward error of $s$-step GMRES with reorthogonalized block classical Gram–Schmidt (BCGSI+)
In this subsection, we consider a different and more complicated situation, for example BCGSI+, which means that we cannot obtain (55) with an exactly orthonormal matrix $\tilde{V}$. According to Lemma 3, we need to prove both (13) and (26). We provide the properties of BCGSI+ in Appendix A.

### v3

@@ Abstract
Communication, i.e., data movement, is a critical bottleneck for the performance of classical Krylov subspace method solvers on modern computer architectures. Variants of these methods which avoid communication have been introduced, which, while equivalent in exact arithmetic, can be unstable in finite precision. In this work, we address the backward stability of $s$-step GMRES, also known as communication-avoiding GMRES. Building upon the “modular framework” proposed in [A. Buttari, N. J. Higham, T. Mary, & B. Vieublé. Preprint in 2024.], we present an improved framework for simplifying the analysis of $s$-step GMRES, which includes standard GMRES ($s=1$) as a special case, by isolating the effects of rounding errors in the QR factorization and the solution of the least squares problem. The key advantage of this new framework is that it is evident how the orthogonalization method affects the backward error, and it is not necessary to re-evaluate anything other than the orthogonalization itself when modifying the orthogonalization used in GMRES. Using this framework, we analyze $s$-step GMRES with popular block orthogonalization methods: block modified Gram–Schmidt and reorthogonalized block classical Gram–Schmidt algorithms.
@@ 1 Introduction
The standard GMRES algorithm creates and orthogonalizes the Krylov basis one vector at a time, predominantly making use of BLAS 2 operations. On modern computer architectures, the performance of classical iterative solvers is heavily dominated by communication, i.e., data movement and synchronization. This motivated the introduction of $s$-step (also called communication-avoiding) variants of the GMRES algorithm, which can reduce asymptotic latency costs by a factor of $s$ and can take advantage of BLAS 3 operations. In [15, Section 3.6], it is demonstrated that CA-GMRES provides a speedup ranging from 1.31×1.31$\times$1.31 × to 4.3×4.3$\times$4.3 × compared to the standard GMRES on an Intel platform with 8 cores. The authors in [23] show that CA-GMRES can achieve speedups of up to 2×2$\times$2 × over the standard GMRES on one or multiple GPUs. Particularly, the orthogonalization phase in CA-GMRES can reach up to 4.16×4.16$\times$4.16 × speedup compared to the same phase in the standard GMRES when $s=10$ is used.
@@ 3 An abstract framework for backward stability of $s$-step GMRES
Notice that the first column of the $Q$-factor associated with $\begin{bmatrix}\phi\widehat{r}&\widehat{W}_{1:(i+1)s}\end{bmatrix}$, specifically $V_{1:(i+1)s+1}$, is determined by Line 2 in Algorithm 1, and this computation is column-wise backward stable. Thus, in GMRES, the orthogonalization method employed in Step 2 can be viewed as being applied to the matrix $\begin{bmatrix}\phi\widehat{r}&\widehat{W}_{1:(i+1)s}\end{bmatrix}$ for any positive scalar $\phi$. We assume the backward stability of the block orthogonalization method used in Step 2; according to [11], most commonly-used block orthogonalization schemes satisfy this property, i.e., for any $\phi>0$,
@@
From standard rounding error analysis and [14, Equation (3.12)], we have $\varepsilon_{r}$, $\varepsilon_{AZ}$, $\varepsilon_{mv}=O(\bm{u})$. In [5, Equation (A.7)], the authors have given $\varepsilon_{ls}=O(\bm{u})$, which comes from applying Givens QR factorization to the upper Hessenberg matrix $\hat{R}_{1:p+1,2:p+1}$ and solving the triangular system as depicted in Step 3.
@@
According to [20, Theorem 2.4], the residual of the least squares problem in Step 3, i.e., $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$, can be bounded by $\sigma_{\min}\bigg{(}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\bigg{)}$ and $\Big{\|}\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}\Big{\rVert}_{\mathsf{F}}$. Therefore, establishing a bound for the residual $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$ requires determining the relationship between the result and the assumptions (13) and (14), which amounts to bounding
@@
Lemma 1 gives bounds on the smallest and largest singular values of the matrices $\begin{bmatrix}\phi\widehat{R}_{1:p+1,1}&\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}\end{bmatrix}$ and $\widehat{R}_{1:p+1,2:p+1}D_{1:p}^{-1}$, which will be required for the proof of Lemma 3. We are now prepared to bound the residual of the least squares problem in Step 3, i.e., $\bigg{\|}\widehat{R}_{1:p+1}\begin{bmatrix}1\\ -\widehat{y}^{(k)}\end{bmatrix}\bigg{\rVert}$, using [20, Theorem 2.4].
@@
We follow the approach from [5, 20] and employ [20, Theorem 2.4] to estimate the residual of the least squares problem. Let $\tilde{R}_{1:p+1}=\widehat{R}_{1:p+1}+\Delta R_{1:p+1}$. Then using [20, Theorem 2.4], the residual of the least squares problem can be estimated as, for any $\phi>0$,
@@ 4.1 Backward error of $s$-step GMRES with block Householder QR and block modified Gram–Schmidt (BMGS)
From [14, Theorem 19.4] and [24], it is easy to see that
@@
Then we consider the modified Gram–Schmidt (MGS) algorithm. [14, Theorem 19.13] and [4] showed (55) can be satisfied by establishing the equivalence between the MGS algorithm and the Householder QR algorithm. Similarly, it has been proved in [2] that applying the BMGS algorithm with Householder QR as the intra-block orthogonalization routine to $X$ is equivalent to applying the block Householder QR algorithm to $\begin{bmatrix}0_{n}\\ X\end{bmatrix}$ both mathematically and numerically. Then we obtain that (55) holds also for the BMGS algorithm.
@@
holds, as established in [17, Theorem 4.1]. Moreover, the contrapositive indicates that (13) is satisfied with $\omega_{\kappa(r,W)}=O(\bm{u})$ when there exists $p\leq n$ such that $\bigl{\|}\widehat{V}_{1:p+1}^{\top}\widehat{V}_{1:p+1}-I\bigr{\rVert}_{\mathsf{F}}>1/2$ but $\bigl{\|}\widehat{V}_{1:p}^{\top}\widehat{V}_{1:p}-I\bigr{\rVert}_{\mathsf{F}}\leq 1/2$.
@@ 4.2 Backward error of $s$-step GMRES with reorthogonalized block classical Gram–Schmidt (BCGSI+)
In this subsection, we examine another scenario, such as BCGSI+. As derived from [9, Corollary 1], it is evident that we can achieve (55) using an exactly orthonormal matrix $\tilde{V}$ for BCGSI+. Nevertheless, in this case, we also offer a more complicated proof under the assumption that obtaining (55) with an exactly orthonormal matrix $\tilde{V}$ is not feasible, as detailed in Appendix A. This analysis may shed light on orthogonalization methods that indeed cannot accomplish (55) with an exactly orthonormal matrix $\tilde{V}$. According to Lemma 3, we need to prove both (13) and (26). We provide the properties of BCGSI+ in Appendix A.